OpenStudy (anonymous):

How many customers does a portrait studio serve in a week if the studio averages 5 customers each weekday and averages twice as many customers on each weekend day?

OpenStudy (anonymous):

customers on weekdays: 5*5=25 customers on weekends: 5*2*2=20 25+20=45

OpenStudy (anonymous):

\[(5*5)for the week days, +(2*10) for the week end\]

OpenStudy (anonymous):

thankyou guys and gals